# supplement

### v1

@@ Abstract
Cultural evolution theory suggests that prestige bias–whereby individuals preferentially learn from prestigious figures–has played a key role in human ecological success. However, its impact within online environments remains unclear, particularly regarding whether reposts by prestigious individuals amplify diffusion more effectively than reposts by non-influential users. Here, we analyzed over 55 million posts and 520 million reposts on Twitter (currently X) to examine whether users with high influence scores (hg-index) more effectively amplified the reach of others’ content. Our findings indicate that posts shared by influencers were more likely to be further shared compared to those shared by non-influencers. This effect persisted over time, especially in viral posts. Moreover, a small group of highly influential users accounted for approximately half of the information flow within repost cascades. These findings demonstrate a prestige bias in information diffusion within digital society, suggesting that cognitive biases shape content spread through reposting.
@@ 1 Introduction
In the digital age, social media influencers have emerged as authoritative figures with significantly greater information diffusion power than ordinary users [1, 2]. The impact of influencers extends beyond the online world, as evidenced by the emergence of influencer-based viral marketing strategies [3, 4, 5] and legal disputes concerning reposts by an influencer on Twitter (currently X) [6]. Influencers are believed to significantly impact the distribution of information in online communities due to their voice and follower count. In recent years, there have various approaches to addressing problems related to online communication, such as the spread of misinformation or homogeneity of received information (so-called echo-chambers and filter bubbles). Research has shown that influencers, particularly experts in specific areas, can contribute to correcting misinformation in online communities [7]. Accordingly, a continued focus on influencers remains crucial in this context.
@@
Previous research examining influencers’ role in information diffusion has focused on their capacity as source spreaders; i,e. their ability to generate and spread original content [1, 8, 2, 9]. However, a key feature of modern social networks comprises repost or share functions, which allow users to reproduce others’ posts; such functionality is a major contributor to information diffusion. Acerbi [10] pointed out that sharing content online fundamentally differs from oral retelling, as it does not require memorization or reproduction of the content. This facilitates and potentially accelerates the spread of information in online environments, contributing to viral-like diffusion patterns. Some researchers have examined influencers as brokers of information, analyzing their role in facilitating the spread of content created by others [11, 12, 9]. Despite these research advances, significant gaps remain. Among these is the question of whether influencers who excel at spreading self-generated content can also amplify other-generated content through reposts. That is, are recipients more likely to share a post simply because an influencer has reposted it? Thus, the present study aims to investigate whether the diffusion capabilities of influencers extend beyond their own content, potentially affecting the propagation of information posted by others. Specifically, this research seeks to explore how these influencers impact the preferences and engagement behaviors of users who receive shared content.
@@
If recipients are indeed more likely to share content from influencers, this would suggest the presence of cognitive biases. In evolutionary anthropology and cultural psychology, researchers have long studied context-dependent biases in information acquisition, where individuals selectively adopt information based on social context rather than content alone. Among these, prestige bias is the tendency to learn from socially successful or prestigious individuals [13, 14, 15]; it is considered an adaptive trait from a historical perspective. This concept aligns with established theories such as the theory of communication (e.g., two-step flow[16] and diffusion of innovations[17]), which emphasize the role of opinion leaders and early adopters in information spread. Recent studies on social media diffusion patterns [18] have indicated that most sharing occurs within close proximity to the original poster, highlighting the importance of early-stage diffusion and influential users. Furthermore, research has shown influencers’s potential to correct misinformation [7]. Building on these theories and recent findings, this study focuses on the early stages of information diffusion. We aim to track and analyze how influence changes over time, examining the effect of prestige bias in the digital contexts characterizing our information-rich society.
@@
When applied to social media, the concept of prestige bias raises important questions: do influencers function as prestigious models whose reposts are more likely to be further reposted than those of non-influencers? Clearly, influencers’ original content is appealing to their audience of active followers; however, it remains uncertain whether they maintain a similar influence when merely reposting others’ information. Understanding if there is a cognitive tendency to prefer any information associated with prestigious individuals, whether original or reposted, would provide useful insights about how information recipients are influenced. Our research thus aims to investigate whether having a proven ability to spread original content (our definition of ”influencer”) translates into an enhanced ability to spread others’ content through reposts. Importantly, this study does not focus on the number of followers an influencer has; rather, it examines individual cognition and behavior. Specifically, it examines whether individuals are more likely to spread information they receive from influencers compared to non-influencers. Our investigation of the early stages of information diffusion allows us to explore the nuanced dynamics of information spread in social media, building insights about the cognitive processes underlying user behavior (content sharing) beyond simple metrics.
@@
Figure 1: (a) Illustration of a repost cascade. The blue speech bubble represents an original post, while the green repost icons represent subsequent reposts. The arrows indicate the direction of information flow. (b) Illustration of primary spread, where an original post (blue speech bubble) is directly shared with multiple users. (c) Illustration of secondary spread, where a user reposts (green icon) content from another user, which is then further shared.
@@
To test our hypothesis about the existence of prestige bias in online communities, we need to precisely track how information flows to subsequent recipients via reposts; this requires the introduction of several key concepts. First, we introduce the concept of a repost cascade, which builds upon the established notion of information cascades [19, 20, 2, 18, 21]. Figure 1 illustrates this concept and shows two key types of information spread that we distinguish in our study. This framework allows us to identify who received information from whom. As shown in Figure 1(a), a repost cascade represents the chain of reposts originating from an original post, capturing the cascading structure of information diffusion on social networks. The blue speech icon (speech bubble) represents an original post, while the green repost icons represent subsequent reposts, with arrows indicating the direction of information flow over time. Next, we distinguish between two types of information spread: primary spread, which is the diffusion of a user’s original post, and secondary spread, which is the diffusion of a user’s repost. Figure 1(b) depicts primary spread, where an original post (blue speech bubble icon) from a source spreader is directly shared with multiple users, including followers and followers’ followers. This process represents the propagation of original content through the network. Secondary spread, shown in Figure 1(c), represents the phenomenon whereby content is reposted by a user (indicated by the green repost icon) and then further spread among their followers. This secondary process captures how information continues to spread beyond its initial audience through the actions of intermediary users. This distinction is crucial for our analysis as it allows us to differentiate between a user’s ability to spread their own content (primary spread) and their capacity to amplify others’ messages (secondary spread). By focusing on these two types of spread, we can investigate whether accounts with powerful primary spread capabilities (influencers, which are quantified by the hg-index[22] in our study) also have powerful secondary spread capabilities.
@@
Figure 2: (a) Virtual timeline: A representation of posts shared by users with different influence levels. (b) Repost cascades: Illustration of how reposts are tracked in cascades. If the edge linking users X and Y appears in the post’s cascade, it is counted as having been reposted. (c) Cascading repost probability (CRP): An example calculation of the CRP for a reposted user with three followers, where two of them reposted, resulting in a CRP of $2/3\approx 0.66$.
@@
Furthermore, to track who views and engages with reposts from specific users, we construct a virtual timeline from sampled reposts and follower-followee relationships, as illustrated in Figure 2(a). This approach allows us to simulate the followers’ exposure to reposts and observe their subsequent engagement. Lastly, using this approach, we introduce the concept of cascading repost probability (CRP) to measure the efficiency of information spread (Figure 2(b,c)). We consider a repost on the virtual timeline as having been further shared if it appears in a repost cascade, allowing us to calculate the probability of a user’s secondary spread continuing (Figure 2(b)). The CRP quantifies the likelihood of a repost being further shared, enabling us to evaluate the efficiency of repost diffusion among users with varying levels of influence; the CRP thus measures a user’s influence in the context of secondary spread. (Details are shown in Figure 2(c).) We calculates the CRP by aggregating view and repost counts across all users’ timelines. By comparing the CRPs of influencers and non-influencers, we can assess whether prestige bias manifests in online environments. If influencers consistently demonstrate higher CRP values, it would suggest that their status enhances their ability to spread information, even when that information originates from others.
@@
These concepts (i.e., repost cascade, secondary spread, virtual timeline, and CRP) provide us with a framework to quantify and analyze users’ influence in terms of secondary information spread. Our methodology focuses on the dynamics of secondary spread and how it relates to users’ level of influence, particularly comparing influencers and non-influencers. Based on this framework, we hypothesize that influencers have higher CRP values compared to non-influencers; these values remain consistently higher over time, though they are especially high in the early stages after the original post has been posted. If supported, this would provide evidence for the effect of prestige bias in online information diffusion. Importantly, while the CRP represents the efficiency of reposts in secondary spread, it does not capture the actual scale of distribution. Therefore, to further support our hypothesis, we also identify the proportion of the actual distribution of reposts in secondary spreads that can be attributed to influencers. This comprehensive approach allows us to examine not only the efficiency of influencers’ information spread, but also their overall impact on information diffusion in online communities.
@@
We thus aim to provide a nuanced understanding of how user status and influence shape the dynamics of information spread in digital environments. Our study thus contributes to the broader understanding of information diffusion in online social networks, the role of influencers in this process, and the potential presence and impact of prestige bias in digital contexts. The results of this investigation have implications for our understanding of online information dynamics, influencer marketing strategies, and the design of social media platforms.
@@ 2.1 Data Collection
Following established methodologies in previous studies [18, 21], we constructed repost cascades based on one month of Japanese-language posts sampled from Twitter (currently X) and their associated follower-followee relationships. The dataset comprised 55,882,528 source posts, 520,048,995 reposts, and 14,910,772 unique users. For detailed information on the data collection process and the rationale behind choosing Japanese-language posts, please refer to the 4.2 in Methods section.
@@ 2.2 User Influence Distribution and Source Post Popularity
To categorize users based on their influence as source spreaders, we employed the hg-index, which quantifies a user’s ability to consistently generate and spread original content. This metric extends the h-index [23], which is commonly used to measure scientific productivity, by incorporating additional factors to provide a more comprehensive measure of a user’s primary spread capability (details in Methods). Using the distribution of hg-index scores, we classified users into six influence categories through quantile binning: very high (top 1%), high (top 1-5%), upper-mid (top 5-10%), mid (top 10-30%), lower-mid (top 30-50%), and low (bottom 50%), where each category excludes its upper threshold. Table 1 presents a comprehensive overview of influence scores and the popularity of users’ original posts as measured by repost statistics for each user category. As shown in Table 1, influencers classified by the hg-index as having very high influence are highly effective source spreaders.
Table 1: Influence scores as source spreaders and popularity of original posts by user category. User influence, No. of users, Avg. hg-index, Avg. reposts (mean, average number of reposts per post for each user category.), Freq. of reposts (maximum, the highest number of reposts received by a single post within each category.)
User influence No. of users Avg. hg-index Avg. reposts Freq. of reposts
@@ 2.3 Analyzing Prestige Bias in Secondary Spread
We explain how the chain of secondary spread(i.e, the impact of user influence on information diffusion) changes for each category. The chain of secondary spread is measured by the CRP. This is calculated by determining whether reposts flowing in the virtual timeline created for each user are reposted by that user before aggregating the results for all users. To determine whether a repost occured, we use repost cascades. When user Y’s repost appears in user X’s virtual timeline, we check the post’s repost cascade to confirm if X subsequently shared that content. If such a repost by X is found in the cascade, we consider the content to have been shared. Please refer to Figure 2 (b) for details. Since we are interested in the initial CRP and how it changes, we focus particularly on the first 6 hours and the temporal dynamics.
@@
Figure 3: Analysis of cascading repost probability (CRP). (a) CRPs by post popularity for different user influence categories, 6 hours after posting. (b) Temporal dynamics of CRPs over 24 hours for posts with $\geq$ 5000 reposts. In both panels, very high influence users consistently show higher CRP values across all popularity thresholds except $<1000$; this advantage is maintained over time.
@@
Figure 3(a) shows the CRP within the first 6 hours after posting, categorized by the minimum number of reposts ($<$ 1000, $\geq$ 1000, $\geq$ 5000, $\geq$ 10000) and user influence categories. The results demonstrate that users with very high influence consistently exhibit higher CRP values for posts with a high number of reposts ($\geq$ 1000), indicating a higher ability to propagate information, even when they are not the original source. Notably, the difference in CRP between influence categories becomes more pronounced as the minimum number of reposts increases, suggesting that the impact of user influence is particularly strong for highly popular content.
@@
To examine the temporal dynamics of this effect, we analyzed the CRP over a 24 hour period for cascades with $\geq$ 5000 reposts, as shown in Figure 3(b). The graph displays the CRP for different influence categories over 24 hours. While CRP generally decreases over time for all categories, users with very high influence maintain substantially higher CRP throughout the observation period, further emphasizing their sustained impact on information diffusion. These findings consistently suggest that user prestige enhances the perceived value or interest of shared content, increasing the likelihood of further diffusion.
Our analysis thus provides strong evidence for the role of prestige bias in information diffusion in online social networks, supporting our hypothesis. Users with higher influence, measured by the hg-index, show greater ability to propagate information, even when it was originally created by someone else. This aligns with the theoretical framework of prestige bias proposed by Henrich and Gil-White [13] and Jiménez and Mesoudi [14], extending this cognitive tendency to digital environments. Notably, this effect is particularly pronounced for posts with high popularity ($\geq$ 1000 reposts) and persists throughout the initial diffusion stages. The CRP for influential users remains consistently higher over time, indicating their status enhances their information spread capability. However, prestige bias appears to more strongly affect popular content ($\geq$ 1000 reposts), suggesting its influence varies across content types. These findings have significant implications for understanding information spread dynamics in social media and influential users’ role in shaping online discourse. The nuanced effect of prestige bias we observed highlights the critical importance of user status in information diffusion, while also indicating that content characteristics play a role.

### v3

# Prestige bias drives the viral spread of content reposted by influencers in online communities

@@ Abstract
Cultural evolution theory suggests that prestige bias—whereby individuals preferentially learn from prestigious figures—has played a key role in human ecological success. However, its impact within online environments remains unclear, particularly with respect to whether reposts by prestigious individuals amplify diffusion more effectively than reposts by noninfluential users. We analyzed over 55 million posts and 520 million reposts on Twitter (currently X) to examine whether users with high influence scores (hg indices) more effectively amplified the reach of others’ content. Our findings indicate that posts shared by influencers are more likely to be further shared than those shared by non-influencers. This effect persisted over time, especially in viral posts. Moreover, a small group of highly influential users accounted for approximately half of the information flow within repost cascades. These findings demonstrate a prestige bias in information diffusion within the digital society, suggesting that cognitive biases shape content spread through reposting.
@@ 1 Introduction
In the digital age, social media influencers have emerged as authoritative figures with significantly greater information diffusion power than ordinary users do [1, 2]. The impact of influencers extends beyond the online world, as evidenced by the emergence of influencer-based viral marketing strategies [3, 4, 5] and legal disputes concerning reposts by an influencer on Twitter (currently X) [6]. Influencers are believed to significantly impact the distribution of information in online communities because of their voice and follower count. In recent years, various approaches have been used to address problems related to online communication, such as the spread of false information or homogeneity of received information (so-called echo chambers and filter bubbles). However, focusing on influences remains crucial in these contexts. Research has shown that influencers, particularly experts in specific areas, can contribute to correcting false information in online communities [7]. Accordingly, a continued focus on influencers remains crucial in this context.
@@
Previous research examining influencers’ role in information diffusion has focused on their capacity as source spreaders; i.e., their ability to generate and spread original content [1, 8, 2, 9]. However, a key feature of modern social networks comprises repost or share functions, which allow users to reproduce others’ posts; such functionality is a major contributor to information diffusion. Acerbi [10] noted that sharing content online fundamentally differs from oral retelling, as it does not require memorization or reproduction of the content. This facilitates and potentially accelerates the spread of information in online environments, contributing to viral-like diffusion patterns. Some researchers have examined influencers as brokers of information, analyzing their role in facilitating the spread of content created by others [11, 12, 9]. Despite these research advances, significant gaps remain. Among these is the question of whether influencers who excel at spreading self-generated content can also amplify other-generated content through reposts. That is, are recipients more likely to share a post simply because an influencer has reposted it? Thus, the present study aims to investigate whether the diffusion capabilities of influencers extend beyond their own content, potentially affecting the propagation of information posted by others. Specifically, this research aims to explore how these influencers impact the preferences and engagement behaviors of users who receive shared content.
@@
If recipients are indeed more likely to share content from influencers, this would suggest the presence of cognitive biases. In evolutionary anthropology and cultural psychology, researchers have long studied context-dependent biases in information acquisition, where individuals selectively adopt information on the basis of social context rather than content alone. Among these, prestige bias is the tendency to learn from socially successful or prestigious individuals [13, 14, 15]; it is considered an adaptive trait from a historical perspective. This concept aligns with established theories such as the theory of communication (e.g., two-step flow[16] and diffusion of innovations[17]), which emphasize the role of opinion leaders and early adopters in information spread. Recent studies on social media diffusion patterns [18] have indicated that most sharing occurs within close proximity to the original poster, highlighting the importance of early-stage diffusion and influential users. Furthermore, research has shown influencers’ potential to correct false information [7]. Building on these theories and recent findings, this study focuses on the early stages of information diffusion. We aim to track and analyze how influence changes over time, examining the effect of prestige bias in the digital contexts characterizing our information-rich society.
@@
When applied to social media, the concept of prestige bias raises important questions: do influencers function as prestigious models whose reposts are more likely to be further reposted than noninfluencers? Clearly, influencers’ original content appeals to their audience of active followers; however, it remains uncertain whether influencers maintain a similar influence when merely reposting others’ information. Understanding whether there is a cognitive tendency to prefer any information associated with prestigious individuals, whether original or reposted, would provide useful insights into how information recipients are influenced. Our research thus aims to investigate whether having a proven ability to spread original content (our definition of ”influencer”) translates into an enhanced ability to spread others’ content through reposts. Importantly, this study does not focus on the number of followers an influencer has; rather, it examines individual cognition and behavior. Specifically, it examines whether individuals are more likely to spread information they receive from influencers than from noninfluencers. Our investigation of the early stages of information diffusion allows us to explore the nuanced dynamics of information spread in social media, building insights into the cognitive processes underlying user behavior (content sharing) beyond simple metrics.
@@
Figure 1: (a) Illustration of a repost cascade. The blue speech bubble represents an original post, whereas the green repost icons represent subsequent reposts. The arrows indicate the direction of information flow. (b) Illustration of primary spread, where an original post (blue speech bubble) is directly shared with multiple users. (c) Illustration of secondary spread, where a user reposts (green icon) content from another user, which is then further shared.
@@
To test our hypothesis about the existence of prestige bias in online communities, we need to precisely track how information flows to subsequent recipients via reposts; this requires the introduction of several key concepts. First, we introduce the concept of a repost cascade, which builds upon the established notion of information cascades [19, 20, 2, 18, 21]. Figure 1 illustrates this concept and shows two key types of information spread that we distinguish in our study. This framework allows us to identify who received information from whom. As shown in Figure 1(a), a repost cascade represents the chain of reposts originating from an original post, capturing the cascading structure of information diffusion on social networks. The blue speech icon (speech bubble) represents an original post, whereas the green repost icons represent subsequent reposts, with arrows indicating the direction of information flow over time. Next, we distinguish between two types of information spread: primary spread, which is the diffusion of a user’s original post, and secondary spread, which is the diffusion of a user’s repost. Figure 1(b) depicts the primary spread, where an original post (blue speech bubble icon) from a source spreader is directly shared with multiple users, including followers and followers’ followers. This process represents the propagation of original content through the network. The secondary spread, shown in Figure 1(c), represents the phenomenon whereby content is reposted by a user (indicated by the green repost icon) and then further spread among their followers. This secondary process captures how information continues to spread beyond its initial audience through the actions of intermediary users. This distinction is crucial for our analysis, as it allows us to differentiate between a user’s ability to spread their own content (primary spread) and their capacity to amplify others’ messages (secondary spread). By focusing on these two types of spread, we can investigate whether accounts with powerful primary spread capabilities (influencers, which are quantified by the hg index[22] in our study) also have powerful secondary spread capabilities.
@@
Figure 2: (a) Virtual timeline: A representation of posts shared by users with different influence levels. (b) Repost cascades: Illustration of how reposts are tracked in cascades. If the edge linking users X and Y appears in the post’s cascade, it is counted as having been reposted. (c) Cascading repost probability (CRP): An example calculation of the CRP for a reposted user with three followers, where two of them repost, resulting in a CRP of $2/3\approx 0.66$.
@@
Furthermore, to track who views and engages with reposts from specific users, we construct a virtual timeline from the sampled reposts and follower‒followee relationships, as illustrated in Figure 2(a). This approach allows us to simulate followers’ exposure to reposts and observe their subsequent engagement. Finally, via this approach, we introduce the concept of cascading repost probability (CRP) to measure the efficiency of information spread (Figure 2(b,c)). We consider a repost on the virtual timeline as having been further shared if it appears in a repost cascade, allowing us to calculate the probability of a user’s secondary spread continuing (Figure 2(b)). The CRP quantifies the likelihood of a repost being further shared, enabling us to evaluate the efficiency of repost diffusion among users with varying levels of influence; the CRP thus measures a user’s influence in the context of secondary spread. (Details are shown in Figure 2(c).)
We calculate the CRP by aggregating the view and repost counts across all the users’ timelines. By comparing the CRPs of influencers and noninfluencers, we can assess whether prestige bias manifests in online environments. If influencers consistently demonstrate higher CRP values, it suggests that their status enhances their ability to spread information, even when that information originates from others.
@@
These concepts (i.e., repost cascade, secondary spread, virtual timeline, and CRP) provide us with a framework to quantify and analyze users’ influence in terms of secondary information spread. Our methodology focuses on the dynamics of secondary spread and how it relates to users’ level of influence, particularly when comparing influencers and noninfluencers. On the basis of this framework, we hypothesize that influencers have higher CRP values than noninfluencers; these values remain consistently higher over time, although they are especially high in the early stages after the original post has been posted. Moreover, we also analyzed this hypothesis with a mixed-effects logistic regression that controlled for post-level and user-level heterogeneity as random effects. If supported, this would provide evidence for the effect of prestige bias in online information diffusion. Notably, while the CRP represents the efficiency of reposts in the secondary spread, it does not capture the actual scale of distribution and indirect influence. Therefore, to further support our hypothesis, we also identified the proportion of the actual distribution of reposts in secondary spreads attributed to influencers and analyzed the impact on max depth and structural virality of a cascade from first-reposted users. This comprehensive approach allows us to examine the efficiency of influencers’ information spread and its overall impact on information diffusion in online communities.
@@
We thus aim to provide insight into how user status and influence shape the dynamics of information spread in digital environments. Our study thus contributes to the broader understanding of information diffusion in online social networks, the role of influencers in this process, and the potential presence and impact of prestige bias in digital contexts. The results of this investigation have implications for our understanding of online information dynamics, influencer marketing strategies, and the design of social media platforms.
@@ 2.1 Data Collection
Following established methodologies in previous studies [18, 21], we constructed repost cascades on the basis of one month of Japanese-language posts sampled from Twitter (currently X) and their associated follower‒followee relationships. The dataset comprises 55,882,528 source posts, 520,048,995 reposts, and 14,910,772 unique users. For detailed information on the data collection process and the rationale behind choosing Japanese-language posts, please refer to the subsection 4.2 in the Methods.
@@ 2.2 User Influence Distribution and Source Post Popularity
To categorize users by their influence as source spreaders, we employed the hg index, which quantifies a user’s ability to consistently generate and spread original content. This metric extends the h index [23], which is commonly used to measure scientific productivity, by incorporating additional factors. Thus, it offers a more comprehensive measure of a user’s primary spread capability (see details in Methods). On the basis of the distribution of the hg index scores, we classified users into six influence categories through quantile binning: very high (top 1%), high (top 1–5%), upper-mid (top 5–10%), mid (top 10–30%), lower-mid (top 30–50%), and low (bottom 50%). Each category excludes its upper threshold. Table 1 summarizes these user categories, showing the number of users, their average hg index, and the popularity of their original posts. As indicated, users in the very high category exhibit the highest average repost counts as well as the highest maximum number of reposts.
Table 1: Influence scores as source spreaders and popularity of original posts by user category. User influence, No. of users, Avg. hg index, Avg. reposts (mean, average number of reposts per post for each user category). , Freq. of reposts (the maximum number of reposts received by a single post within each category).
User influence No. of users Avg. hg index Avg. reposts Freq. of reposts
@@ 2.3 Analyzing Prestige Bias in Secondary Spread
We next explored how user influence impacts the secondary spread of information—that is, how content diffuses after being reposted by users other than the original poster. We measure this chain of secondary spread via the cascading repost probability (CRP). First, we construct a virtual timeline for each user, tracking when reposts appear. Then, to determine whether a repost occurred, we inspect repost cascades: if user Y’s repost appears in user X’s virtual timeline, we check whether X subsequently shares the content. If so, we count it as having spread by Y. Finally, we aggregate the repost counts by users’ influences. Figure 2(b) illustrates this process. Because we focus on initial diffusion, we aggregate the results over the first 6 h after a post appears.
@@
Figure 3: Analysis of the cascading repost probability (CRP). (a) CRPs by post popularity for different user influence categories 6 h after posting. (b) Temporal dynamics of CRPs over 24 h for posts with $\geq$ 5000 reposts. The x-axis indicates hours since the original post was published. In both panels, very high-influence users consistently show higher CRP values across all popularity thresholds except $<1000$; this advantage is maintained over time.
@@
Figure 3(a) shows the CRP within the first 6 h after posting, categorized by the minimum number of reposts ($<$ 1000, $\geq$ 1000, $\geq$ 5000, $\geq$ 10000) and user influence categories. The results demonstrate that users with very high influence consistently exhibit higher CRP values for posts with a high number of reposts ($\geq$ 1000), indicating a greater ability to propagate information, even when they are not the original source. Notably, the difference in the CRP between influence categories becomes more pronounced as the minimum number of reposts increases, suggesting that the impact of user influence is particularly strong for highly popular content.
@@
To examine the temporal dynamics of this effect, we analyzed the CRP over a 24-h period after the original post time for cascades with $\geq$ 5000 reposts, as shown in Figure 3(b). The x-axis in Figure 3(b) represents hours elapsed since the original post. The graph displays the CRP for different influence categories over 24 h. Although the CRP generally decreases over time for all categories, users with very high influence maintain substantially higher CRP throughout the observation period, further emphasizing their sustained impact on information diffusion.
These findings collectively demonstrate that user prestige enhances the perceived value or interest of shared content, thereby increasing its likelihood of further diffusion. Our analysis shows that prestige bias plays a pivotal role in online social networks, which is consistent with our initial hypothesis. In particular, users with greater influence, as measured by the hg index, exhibit a greater capacity to propagate information, even if it originates from others.
Figure 4: Cascading repost probability (CRP) analysis using English-language posts collected on January 1, 2015. The left panel (a) shows the results restricted to 30 min, and the right panel (b) compares 30 min, 1 h, and 3 h.
Finally, to verify the robustness of our results, we examined a dataset of English-language posts collected on January 1, 2015 (see Supplementary Material S1). We used a sample of one million users from December 2014, provided by [24][25], and collected posts from X (formally Twitter)’s stream over 24 h on January 1. Because this time window is shorter than in our main analysis, we focused on shorter intervals. In Figure 4, panel (a) shows the CRP during the first 30 min, and panel (b) compares the results after 30 min, 1 h, and 3 h. Despite this limited period, a consistent pattern was found: very high-influence users still exhibit high CRP. Interestingly, in this dataset, the low category outperforms mid. This is possibly due to misclassifications, as the hg index is calculated from a single day’s data (some influential users may not have posted on that day). Moreover, this English-language dataset captures only one day of user behavior, so it may not fully capture longer-term diffusion patterns. Despite these minor differences, the analysis overall replicates our main findings in the Japanese dataset: users with greater influence tend to be more effective at propagating content, even under different linguistic and temporal conditions.
2.4 Analysis via Mixed-Effects Logistic Regression
In the analysis presented in subsection 2.2, we did not consider the possibility that influencers select potentially attractive posts that could be reposted by other users. Furthermore, we did not consider the effects of specific posts or users. To exclude these confounding factors, we applied mixed-effects logistic regression accounting for both user-level and post-level heterogeneity. Because the dataset is fairly large, a sampling procedure was used to retain approximately 1% of all positive (reposted) cases along with a twofold number of negative cases. The complete details of the models, dataset, and sampling approach are provided in the Methods section.
In this model, the dependent variable is a binary indicator of whether the user who received a repost subsequently reposts it. The model includes the original reposting user’s influence category (sender_influence), the repost’s hour of the day (repost_hour), and log-transformed and standardized follower counts for both the user who reposted a post (sender_followers_count) and the user who received the repost (user_followers_count). Random intercepts are introduced for each original tweet (source_tweet_id) and the topic of the user profiles (user_topic), thereby controlling for differences in content appeal and individual interests.
Table 2: Fixed-effects estimates from the mixed-effects logistic regression (sender_influence baseline = low).
Predictor $\beta$ value Std. Error $z$ value $p$ value
(Intercept) -0.927 0.137 -6.777 $<$0.001
sender_influence1 (lower-mid) 0.148 0.006 23.085 $<$0.001
sender_influence2 (mid) 0.246 0.007 37.489 $<$0.001
sender_influence3 (upper-mid) 0.287 0.006 45.094 $<$0.001
sender_influence4 (high) 0.387 0.006 66.123 $<$0.001
sender_influence5 (very high) 0.599 0.006 99.480 $<$0.001
repost_hour: morning -0.027 0.003 -8.227 $<$0.001
repost_hour: noon 0.034 0.003 9.813 $<$0.001
repost_hour: night 0.031 0.003 10.010 $<$0.001
repost_hour: midnight -0.018 0.004 -4.626 $<$0.001
sender_followers_count -0.165 0.001 -114.126 $<$0.001
user_followers_count -0.568 0.001 -461.838 $<$0.001
Table 2 presents the fixed-effects estimates for the model, with low influence serving as the baseline. The influence categories all have positive coefficients relative to this baseline, and the very high category has a substantial positive coefficient. Although these estimates cannot be used to infer absolute repost probabilities because of our sampling ratio, the relative differences between categories are meaningful. The effect of very high ($\beta=0.599$) suggests that top-tier influencers can markedly amplify the likelihood of secondary spreading, which is consistent with a prestige bias in online diffusion. The other predictors show comparatively modest but significant effects. Variations by hour of day point to diurnal patterns in user engagement, with slight increases in reposting at night and noon. Both sender_followers_count and user_followers_count have negative coefficients, indicating that once the influence category and other factors are considered, having a larger follower base does not necessarily promote further reposting of others’ content.
Furthermore, our analysis indicates that the random intercept for source_tweet_id exhibits a large variance, suggesting that content-specific factors play an important role in online diffusion. However, our main hypothesis focuses on the effect of user influence rather than the content-specific aspects of the posts. Therefore, we do not present an analysis of content-specific factors in the main text. Nevertheless, we present further details of a topic-specific analysis in the Supplementary Material S3 for reference.